OpenStudy (anonymous):

What is the slope of the line passing through the points (0, 4) and (−8, −1)?

OpenStudy (anonymous):

\[y−y_1=m(x−x_1)\] \[(y−y_1)/(x−x_1)=m\] m = slope of line (x,y) and (x1,y1) are the points

OpenStudy (anonymous):

y - y1 = m(x-x1) (y-y1) /(x-x1) = m

OpenStudy (anonymous):

(4-(-1))/(0-(-8))=5/8

OpenStudy (anonymous):

thanks everybody

OpenStudy (anonymous):

anytime. :)